1. Write the algorithm and draw the flowchart. Given three numbers A, B, C. Compute and print out the sum, the product and the average of these values.
1
Expert's answer
2021-10-29T12:17:04-0400


1  Read A
2  Set Sum = A
3  Set Prod = A
4  Read B
5  Set Sum = Sum + B
6  Set Prod = Prod * B
7  Read C
8  Set Sum = Sum + C
9  Set Prod = Prod * C
10 Set Avg = Sum / 3
11 Print Sum
12 Print Prod
13 Print Avg